OpenStudy (anonymous):

Find the slope and y-intercept of each line. 3x - 4y = 12

OpenStudy (anonymous):

re arrange to y=mx+b 4y=-3x-12 divide by 4: y=-3/4x-3 slope=-3/4 y-int.=-3

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

you're welcome